OpenStudy (anonymous):

A spherical model of planet earth has a radius of 3 ft. What is the approximate volume of the model of planet earth? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

The formula to find the volume of a sphere is \(\rm V = \dfrac{4}{3}\pi r^3\) r = radius, so let's plug that in \(\rm V = \dfrac{4}{3}*3.14* 3^3\)

OpenStudy (anonymous):

i think it is 113.04 but dont know how to explain it in hundreds

sammixboo (sammixboo):

Yes it is 113.04 give me a sec

OpenStudy (anonymous):

ok

sammixboo (sammixboo):

You see the HUNDREDTHS placement here? http://barrstudent126.wikispaces.com/file/view/place-value-chart.gif/331158946/342x234/place-value-chart.gif

sammixboo (sammixboo):

Not hundreds, hundredths

OpenStudy (anonymous):

oh thanks

sammixboo (sammixboo):

It is already in hundredths, but I wanted to give you that chart to help :)

sammixboo (sammixboo):

If you need anymore help let me know! :)

OpenStudy (anonymous):

ok i will thank you